Given the confidence interval for a mean of (74.5738,77.4262), from a sample of size 34 with a population standard deviation of σ=3.6, find the following:

Margin of Error(ME)=

Standard Error(SE)=

Zc=

what was the confidence level for this confidence interval?=

(Show your work please)
Answer #1

confidence interval is                          
lower limit =    74.5738                      
upper limit=   77.4262                      
sample mean = (lower limit+upper limit)/2= (   74.5738   +   77.4262   ) / 2 =   76      
margin of error = (upper limit-lower limit)/2= (   77.4262   -   74.5738   ) / 2 =   1.4262      
           Standard Error , SE = σ/√n =   3.6000   / √    34   =   0.6174  

               


margin of error ,E = Z(*σ/√n )
Z (α/2)= E*√n /σ =   1.426   * √    34   /   3.600   =   2.310
                          
α/2 from critical value    2.310   is                  
α/2=   0.0104                      
α=   0.0209                      
                          
confidence level = 1-α=   1-   0.0209   =   0.9791 or 97.91%